# Statistically-Guided Dual-Domain Meta-Learning with Adaptive Multi-Prototype Aggregation for Distributed Fiber Optic Sensing

### v1

@@ 5 Experiments
All experiments are run on an NVIDIA vGPU-32GB platform with batch size fixed at 32. We evaluate on two real-world cross-deployment datasets, OSDG1 and OSDG2, and organize results into: (i) ablations of each component of our framework; (ii) comparisons with conventional deep models; and (iii) evaluations against meta-learning baselines. We report Accuracy, Precision, Recall, and F1.
For meta-learning, we adopt a domain-generalization episodic protocol where support samples are drawn from source deployment(s) (TRAIN) and queries from a held-out target deployment (TEST), with class-balanced query sets. Unless otherwise stated: on OSDG1 we use N=4, K=4, Q=12, 1,000 episodes; on OSDG2 we use N=3, K=3, Q=12, 1,000 episodes; all with seed 222. Episodes are sampled over the class intersection with at least K+QK{+}Q samples per class. Because queries are class-balanced, aggregated Accuracy equals Macro-Recall for the meta-learning rows. All baselines follow the same construction; gradient-based methods use 10 inner steps with inner LR 0.01 at test time, while metric/closed-form methods perform no inner-loop adaptation. None of the methods used target labels; they were purely domain generalizations.
Table 3: Comparison of Deep Learning and Meta-Learning Methods on OSDG1 and OSDG2 Datasets.
@@
Accuracy Precision Recall F1 Accuracy Precision Recall F1
Xgboost(Chen and Guestrin, 2016) 0.1463 0.1201 0.2884 0.1339 0.6440 0.7931 0.6849 0.6433
SVM(Qu et al., 2010) 0.1615 0.0793 0.3213 0.1259 0.6459 0.7915 0.6866 0.6255
KNN(Jia et al., 2019a) 0.4725 0.5043 0.4225 0.3669 0.6384 0.7614 0.6685 0.6165
1DCNN(Wu et al., 2019a) 0.4421 0.2240 0.3986 0.2458 0.6723 0.6316 0.6248 0.6203
2DCNN(Xu et al., 2018) 0.5914 0.6687 0.6520 0.5476 0.6534 0.7828 0.6915 0.6104
LSTM(Sherstinsky, 2020) 0.1524 0.1663 0.3019 0.1759 0.7363 0.6091 0.6551 0.5805
Transformer(Vaswani et al., 2017) 0.1951 0.1997 0.2850 0.1861 0.6459 0.6610 0.6427 0.6224
ProtoNet(Snell et al., 2017) 0.4320 0.2218 0.4320 0.2931 0.7501 0.8087 0.7501 0.7563
MAML(Finn et al., 2017) 0.4817 0.4503 0.4817 0.4358 0.7011 0.7207 0.7011 0.6812
ANIL(Raghu et al., 2019) 0.4262 0.4171 0.4262 0.4020 0.6501 0.6774 0.6501 0.6585
R2D2(Bertinetto et al., 2018) 0.4758 0.4563 0.4758 0.4345 0.7400 0.7323 0.7400 0.7327
FEAT(Ye et al., 2020) 0.4907 0.5143 0.4907 0.4478 0.6786 0.6988 0.6786 0.6526
DUPLE (Ours) 0.6372 0.6347 0.6372 0.6244 0.8195 0.8158 0.8195 0.8167
5.1 Comparative Experiment
To evaluate the effectiveness of our proposed DUPLE framework, we conducted comparative experiments on the OSDG1 and OSDG2 datasets and compared it with several traditional deep learning and meta-learning methods. To ensure comprehensive testing of the model’s performance in handling domain generalization problems, in the OSDG1 dataset, we used all data from one distribution as training data and all data from the other distribution as the test dataset, completely isolating them. The same approach was taken with OSDG2, using data from both distributions for training and data from the other distribution for testing.
Table 3 presents the results of the comparison in terms of four key classification metrics: Accuracy, Precision, Recall, and F1 Score. We include both traditional models (XGBoost, SVM, KNN) as well as deep learning models (1DCNN, 2DCNN, LSTM, Transformer), and meta-learning methods (ProtoNet, MAML, ANIL, R2D2, and FEAT).
We first consider the performance of traditional machine learning methods, which include XGBoost, SVM, and KNN. These models rely on hand-crafted features, such as statistical moments and signal transformations like the Short-Time Fourier Transform (STFT), to capture relevant patterns from the raw data.XGBoost and SVM performed relatively poorly, especially on the OSDG1 dataset, with accuracies below 0.2. This highlights the difficulty of learning robust features from unevenly distributed training data with domain generalization problems. KNN outperformed XGBoost and SVM on the OSDG1 dataset, achieving an accuracy of 0.4725. However, its performance on the OSDG2 dataset was slightly inferior to XGBoost and SVM.
Among the deep learning methods, 2DCNN achieved the highest Accuracy on both datasets, particularly excelling on OSDG1 with an Accuracy of 0.5914. However, the Recall of 2DCNN on OSDG1 was lower than that of the meta-learning models, with a Recall of 0.6520, indicating that while 2DCNN had a reasonable overall performance, it struggled to capture some of the minority classes effectively.1DCNN and LSTM performed relatively poorly, achieving accuracies of 0.4421 and 0.1524 respectively on the OSDG1 dataset, with even lower recall. These models struggle to cope with domain drift and uneven data distribution, which are crucial for the DFOS task with its variable deployment conditions.
In domain generalization tasks, meta-learning methods such as ProtoNet, MAML, ANIL, and R2D2 perform significantly better, especially with limited and unevenly distributed data. ProtoNet, employing a prototype-based approach, achieves accuracies of 0.4320 and 0.7501 on OSDG1 and OSDG2, respectively, with a precision of 0.8087 on the OSDG2 dataset. While ProtoNet performs well on OSDG2, its generalization ability on OSDG1 is weaker, possibly because domain shifts cannot be well captured by individual prototypes for each class.
Another meta-learning method, MAML, outperforms ProtoNet on the OSDG1 dataset, achieving an accuracy of 0.4817 and significantly improved precision. R2D2 also achieves good results, with accuracies of 0.4758 and 0.7400 on the OSDG1 and OSDG2 datasets, respectively. However, none of these methods can match the performance of our proposed framework, DUPLE.
Our DUPLE framework outperforms all other methods on both datasets, achieving a state-of-the-art accuracy of 0.6372 on OSDG1 and 0.8195 on OSDG2. DUPLE also boasts the highest F1 scores on both datasets, further demonstrating its robustness in balancing precision and recall in challenging scenarios.
In summary, our DUPLE framework not only outperforms conventional deep learning models but also surpasses existing meta-learning methods in terms of Accuracy, Recall, and overall F1 Score. This highlights the advantages of our domain-generalization approach, which combines dual-domain features, statistical guidance, and adaptive prototype aggregation for robust and efficient DFOS activity recognition across various deployment configurations.
5.2 Ablation Experiment
In the ablation experiments, we used the same scenario-based protocol and hyperparameters as in the main experiments to evaluate the contributions of the three main modules: the Two-Domain Multi-Prototype Learner (FPM), the Statistically Guided Network (SGN), and the Query-Aware Collaborative Decision Making (CDM). All variants used the same data partitioning, sampling rules, and evaluation settings to ensure reproducibility of the ablation experiments.
Table 4: Ablation Study Results on OSDG1
FPM SGN CDM Acc Precision Recall F1
0.5210 0.5224 0.5210 0.5201
0.5977 0.6192 0.5977 0.5755
0.6172 0.6407 0.6172 0.5966
0.5960 0.6078 0.5960 0.5795
0.6372 0.6347 0.6372 0.6244
Table 5: Ablation Study Results on OSDG2
FPM SGN CDM Acc Precision Recall F1
0.6765 0.7211 0.6765 0.6817
0.7946 0.8064 0.7946 0.7984
0.7990 0.8097 0.7990 0.8024
0.7935 0.8046 0.7935 0.7971
0.8195 0.8158 0.8195 0.8167
We conducted systematic ablation experiments on the FPM, SGN, and CDM modules separately, and the results were consistent across both datasets. First, FPM is the main driver of performance improvement, effectively characterizing intra-class diversity across deployments through its two-domain multi-prototype representation. Adding SGN on top of this provides stable and reproducible small gains because it provides priors on domain importance and prototype sensitivity based on sample statistics, making fusion and matching more robust. Using CDM alone without introducing SGN offers limited improvement, but combining the two creates a query-aware aggregation guided by reliable statistical priors, achieving further marginal benefits on top of FPM. Overall, the complete model achieved double-digit accuracy improvements on both datasets compared to the module-free baseline, with F1 and accuracy improving in tandem: from approximately 0.52 to 0.64 on OSDG1 and from approximately 0.68 to 0.82 on OSDG2. This indicates that our method is not only more accurate but also more stable, significantly mitigating the domain shift problem caused by deployment differences.
In general, the performance patterns on both datasets are consistent. FPM explains most of the performance improvement by using adaptive multi-prototypes in both domains to capture intra-class diversity. SGN provides stable, sample-conditional gains by adjusting domain dependencies and prototype strictness. CDM itself shows little change, but when combined with SGN, it provides additional, repeatable improvements. This suggests that the focus on class prototypes should be guided by statistical priors for each query, rather than applied unguidedly. In summary, FPM provides the main gain, SGN brings stable performance, CDM has limited effect when used alone but is effective when used in conjunction with SGN. The complete model improves performance by about 10 to 15 percentage points on the OSDG1 and OSDG2 datasets compared to the module-free baseline, verifying the key role of statistical prior-guided dual-domain multi-prototype plus query-aware aggregation in cross-deployment generalization.

### v2

@@ 5 Experiments
All experiments were implemented using PyTorch on an NVIDIA vGPU-32GB platform. We evaluate our framework on two real-world cross-deployment datasets, OSDG1 and OSDG2, utilizing a rigorous Leave-One-Deployment-Out (LODO) cross-validation protocol. Specifically, we iteratively designate one deployment as the unseen target domain (TEST) while training on the remaining deployments (TRAIN), reporting the mean Accuracy, Precision, Recall, and F1 Score across all folds (2 folds for OSDG1, 3 folds for OSDG2). This protocol strictly simulates real-world deployment shifts, where the high variance in our reported results reflects the varying degrees of domain discrepancy across different physical environments.
For meta-learning evaluations, we adopt a standard episodic protocol ($N$-way $K$-shot $Q$-query). We define N=4N{=}4 for OSDG1 and N=3N{=}3 for OSDG2, with a fixed query size of Q=12Q{=}12 to ensure class balance. The main comparative experiments are conducted under a 5-shot setting (K=5K{=}5), evaluated over 1,000 randomly sampled episodes to ensure statistical reliability. Consistent with the domain generalization setting, no target domain labels are utilized for model adaptation during inference.
5.1 Comparative Experiment
To rigorously evaluate the efficacy of the proposed DUPLE framework, we conducted comprehensive comparative experiments on the OSDG1 and OSDG2 datasets. We benchmarked our method against a diverse set of baselines, ranging from traditional machine learning algorithms to standard deep learning models, domain generalization (DG) techniques, and state-of-the-art meta-learning frameworks.
Table 7: Per-deployment macro-F1 breakdown under LODO evaluation. Meta-learning models are evaluated at 5-shot setting.
Model OSDG1 (Deployments, %) OSDG2 (Deployments, %)
Wire Mesh Fence Barbed Wire Hangnail Net Railing Wall
GroupDRO $55.62$ $40.83$ $89.05$ $59.84$ $55.56$
ProtoNet $22.73$ $34.44$ $70.98$ $69.86$ $54.77$
FEAT $39.52$ $52.68$ $46.37$ $67.15$ $37.98$
DUPLE (Ours) $60.48$ $65.71$ $95.02$ $81.80$ $51.81$
Table 8: Few-Shot F1 Comparison on OSDG1 and OSDG2 Datasets (in %).
Model OSDG1 (Shots, %) OSDG2 (Shots, %)
1-shot 3-shot 5-shot 10-shot 1-shot 3-shot 5-shot 10-shot
ProtoNet(Snell et al., 2017) $26.20\pm 1.40$ $24.30\pm 2.08$ $24.28\pm 2.19$ $23.54\pm 3.08$ $64.14\pm 9.90$ $65.56\pm 9.83$ $65.20\pm 9.05$ $65.30\pm 8.52$
MAML(Finn et al., 2017) $35.83\pm 10.87$ $35.09\pm 12.51$ $34.65\pm 12.43$ $34.27\pm 12.12$ $61.55\pm 15.00$ $62.33\pm 13.11$ $63.05\pm 13.21$ $63.12\pm 13.49$
ANIL(Raghu et al., 2019) $0.15\pm 6.99 33$$42\pm 10.34 33.$$43\pm 10.19 33.$$3.88\pm 11.25$ $57.92\pm 21.77$ $60.12\pm 21.38$ $60.23\pm 22.78$ $60.18\pm 24.63$
R2D2(Bertinetto et al., 2018) $38.20\pm 16.70$ $37.21\pm 16.03$ $36.87\pm 16.00$ $36.86\pm 15.85$ $54.91\pm 19.59$ $63.67\pm 22.69$ $64.10\pm 23.90$ $64.02\pm 25.15$
FEAT(Ye et al., 2020) $44.28\pm 7.75$ $45.43\pm 10.92$ $6.10\pm 9.31 44$$4.30\pm 9.81$ $48.71\pm 7.62$ $51.25\pm 13.80$ $50.50\pm 15.02$ $49.97\pm 16.65$
DUPLE (Ours) $58.58\pm 2.60$ $62.70\pm 2.98$ $63.09\pm 3.69$ $3.92\pm 3.19 66$$6.43\pm 18.26$ $75.36\pm 21.59$ $76.21\pm 22.14$ $76.02\pm 24.42$
Table 5 summarizes the performance of non-meta-learning approaches. First, traditional machine learning methods such as XGBoost and SVM exhibit limited efficacy in these cross-deployment scenarios. On the OSDG1 dataset, their accuracies remain below 30%, indicating that manual feature engineering is insufficient to capture invariant patterns amidst severe domain shifts.
Second, standard deep learning models demonstrate significant performance variability driven by architecture bias. On OSDG1, the 2DCNN utilizes Time-Frequency (STFT) representations to achieve 42.15% accuracy, which is notably superior to raw-signal-based models like the 1DCNN at 34.64% and the Transformer at only 22.03%. Conversely, the performance trend on OSDG2 shifts dramatically to favor sequence-based models, where the Transformer and LSTM models achieve leading accuracy rates of 68.69% and 67.72%, respectively, significantly outperforming convolutional architectures. This inconsistency highlights that standard supervised learning architectures often overfit to domain-specific feature modalities.
Table 9: Statistical significance (Welch’s $t$-test) against the best meta-learning baseline over 1,000 episodes.
Dataset Best Baseline $\dagger$ DUPLE (Ours) $p$-value Significant?
OSDG1 $49.53\pm 9.07$ (FEAT) $63.07\pm 3.56$ $10^{-4}$ Yes
OSDG2 $66.49\pm 10.77$ (ProtoNet) $77.45\pm 20.35$ $10^{-4}$ Yes
$\dagger$ The best baseline varies by dataset: FEAT for OSDG1, ProtoNet for OSDG2.
Most notably, the Domain Generalization (DG) baselines, GroupDRO and Deep CORAL, demonstrate superior robustness compared to standard Empirical Risk Minimization (ERM) models. To ensure a fair comparison, we implemented these DG algorithms using the same 2DCNN backbone as the standard baseline. By explicitly optimizing for worst-case groups or aligning feature distributions, these methods establish a strong performance baseline. On the challenging OSDG1 dataset, GroupDRO reaches an accuracy of 57.82%, surpassing the standard 2DCNN by a margin of over 15%. Similarly, on OSDG2, GroupDRO and Coral maintain the lead with accuracies of 71.72% and 70.10%, respectively. However, despite their higher average accuracy, these DG methods exhibit high variance, exemplified by the standard deviation of 16.10% observed for GroupDRO on OSDG1. This instability underscores a critical limitation: purely invariant representation learning without test-time adaptation is often insufficient to guarantee consistent performance across highly diverse optical fiber deployment environments.
Tables 6 and 8 present the few-shot classification results. In general, meta-learning baselines exhibit improved generalization capabilities compared to standard deep models, particularly in low-data regimes. On OSDG1, FEAT emerges as the strongest baseline with a 5-shot accuracy of $49.53\%$. In contrast, standard metric-based methods such as ProtoNet perform suboptimally, yielding only $31.55\%$. This underperformance suggests that a unimodal single-prototype representation is inadequate for modeling the complex, multi-modal distribution of fiber signals in this highly shifted scenario. On OSDG2, ProtoNet demonstrates robust performance by achieving $66.49\%$ accuracy, which outperforms more complex architectures like R2D2 and ANIL that reach $64.00\%$ and $63.31\%$ respectively. This indicates that for certain deployment configurations where the domain gap is moderate, the feature space is sufficiently discriminative for Euclidean-distance-based metrics.
The proposed DUPLE framework consistently establishes a new state-of-the-art across both datasets and all shot settings, effectively mitigating the performance gaps observed in baseline methods. Specifically, on the challenging OSDG1 dataset (5-shot), DUPLE achieves an accuracy of 63.07%, surpassing the best-performing meta-learning baseline, FEAT, by a substantial margin of over 13.5%. This significant improvement validates the efficacy of our dual-domain fusion strategy in extracting domain-invariant representations under conditions where single-domain baselines struggle. Similarly, in the OSDG2 scenario, our method attains an accuracy of 77.45%. This performance not only surpasses the strongest meta-learning baseline, ProtoNet, by approximately 11%, but also outperforms the leading domain generalization baseline, GroupDRO (71.72%), by a clear margin of 5.73%.
Notably, we report results as mean $\pm$ standard deviation over LODO folds (OSDG1: 2 deployments; OSDG2: 3 deployments), where the standard deviation reflects cross-deployment heterogeneity rather than random instability. In particular, OSDG2 exhibits larger variability because different held-out deployments pose substantially different difficulty levels. As evidenced by the per-deployment macro-F1 breakdown in Table 7, DUPLE achieves $95.02\%$ on Hangnail net and $81.80\%$ on Railing, but drops to $51.81\%$ on Wall, indicating that Wall constitutes a significantly more challenging deployment scenario. Similar cross-deployment fluctuations are observed for ProtoNet and GroupDRO, confirming that the higher standard deviations primarily arise from real deployment shifts captured by the LODO protocol. Overall, DUPLE’s superior mean performance amid such heterogeneity confirms its robustness under both moderate and severe domain shifts.
To verify that these performance gains are statistically reliable rather than artifacts of random episodic sampling, we conducted Welch’s $t$-tests over 1,000 testing episodes. As detailed in Table 9, we compared DUPLE against the strongest baseline for each dataset. The resulting $p$-values are significantly below $10^{-4}$ in both scenarios, providing strong statistical evidence that DUPLE yields a significant performance improvement over existing state-of-the-art methods.
In conclusion, DUPLE not only surpasses conventional deep models and DG algorithms in generalization capability but also addresses the representational limitations of existing meta-learners. By effectively fusing dual-domain features with adaptive prototype aggregation, it achieves the highest Accuracy and F1 Scores across heterogeneous deployment environments.
5.2 Ablation Experiment and Mechanism Analysis
5.2.1 Ablation Experiment
To systematically dissect the contribution of each component within the DUPLE framework, we conducted ablation studies using the same rigorous Leave-One-Deployment-Out protocol as the main experiments. We evaluated three core modules: the Dual-Domain Multi-Prototype Learner (FPM), the Statistical Guidance Network (SGN), and the Query-Aware Collaborative Decision Making (CDM). The results are summarized in Table 10.
Table 10: Ablation Study of DUPLE Modules (5-Shot). Results in %.
Module OSDG1 (%) OSDG2 (%)
FPM SGN CDM Accuracy Precision Recall F1 Accuracy Precision Recall F1
$49.86\pm 3.17$ $50.22\pm 2.85$ $49.86\pm 3.17$ $49.28\pm 3.86$ $60.03\pm 20.47$ $58.98\pm 22.07$ $60.03\pm 20.47$ $57.35\pm 22.15$
$57.79\pm 2.80$ $59.09\pm 4.01$ $57.79\pm 2.80$ $56.20\pm 1.91$ $66.62\pm 24.60$ $70.81\pm 21.93$ $66.62\pm 24.60$ $64.89\pm 26.18$
$61.07\pm 0.93$ $62.33\pm 2.47$ $61.07\pm 0.93$ $59.88\pm 0.30$ $66.70\pm 24.55$ $70.94\pm 21.63$ $66.70\pm 24.55$ $64.98\pm 26.14$
$60.14\pm 0.75$ $60.85\pm 0.09$ $60.14\pm 0.75$ $59.20\pm 1.75$ $67.53\pm 23.70$ $71.46\pm 20.93$ $67.53\pm 23.70$ $65.67\pm 25.45$
$63.07\pm 3.56$ $64.08\pm 4.35$ $63.07\pm 3.56$ $63.92\pm 3.19$ $77.45\pm 20.35$ $82.87\pm 11.94$ $77.45\pm 20.35$ $76.21\pm 22.14$
We conducted systematic ablation experiments to isolate the contributions of the FPM, SGN, and CDM modules, yielding consistent improvements across both datasets. First, the Dual-Domain Multi-Prototype Learner (FPM) acts as the primary performance driver, effectively characterizing intra-class diversity across deployments; it alone boosts accuracy from the baseline’s $49.86\%$ to $57.79\%$ on OSDG1. Building on this foundation, the Statistically Guided Network (SGN) and Collaborative Decision Making (CDM) modules provide complementary gains. SGN utilizes sample statistics to regularize domain importance, while CDM dynamically refines prototype weights based on the query. Contrary to unguided aggregation, the inclusion of either module yields distinct performance boosts, confirming that both statistical priors and query-aware attention are essential for robust feature alignment under domain shifts.
Overall, the integration of all three modules in the complete DUPLE framework delivers remarkable results. The full model achieves double-digit accuracy improvements compared to the naive baseline, advancing from $49.86\%$ to $63.09\%$ on OSDG1 and from $60.03\%$ to $77.45\%$ on OSDG2. Notably, while the standard deviation on OSDG2 remains high ($\pm 20.35\%$) due to the intrinsic difficulty variances captured by our Leave-One-Deployment-Out protocol, the mean performance consistently surpasses all ablated variants. By combining dual-domain prototypes with statistical and query-based guidance, our approach significantly mitigates the domain shift problem, offering a stable and accurate solution for diverse deployment environments.
5.2.2 Robustness to Data Scarcity
In practical engineering deployments, collecting labeled data for new scenarios is often costly and labor-intensive. To rigorously assess the efficacy of DUPLE in low-data regimes, we conducted experiments with support set sizes ranging from $K=1$ to $10$ on both OSDG1 and OSDG2 datasets. Beyond Accuracy, we also analyzed Precision and F1-Score to assess the trade-off between false alarms and missed detections. Figure 5 presents the results.
Figure 5: Sensitivity analysis of model performance with respect to support set size $K\in\{1,3,5,10\}$. Top Row (OSDG1): (a-c) DUPLE consistently achieves the highest Accuracy, Precision, and F1-Score, demonstrating exceptional robustness even in the extreme low-data regime ($K=1$), while baselines like ProtoNet struggle significantly. Bottom Row (OSDG2): (d-f) DUPLE exhibits superior data efficiency, showing a rapid performance surge from $K=1$ to $K=3$ and establishing a comprehensive lead in all metrics, unlike baselines that plateau at lower performance levels.
On the challenging OSDG1 dataset, DUPLE demonstrates a commanding advantage over all baselines. As illustrated in Figures 5(a) and (c), our method maintains a significant lead in both Accuracy and F1 Score across all $K$ values. Notably, under extreme single-sample conditions ($K=1$), DUPLE achieves an F1 Score exceeding $58\%$, whereas the strongest baseline (FEAT) hovers around $44\%$, and standard metric-based methods like ProtoNet drop below $27\%$. This substantial gap confirms that the SGN module successfully prevents model collapse even with minimal supervision. Furthermore, as depicted in Figure 5(b), DUPLE achieves the highest Precision across the board, indicating its ability to learn robust discriminative features without succumbing to the low-precision pitfalls seen in baseline methods.
For the OSDG2 dataset, the results reveal DUPLE’s superior learning potential. As shown in Figures 5(d) and (f), while DUPLE starts with a moderate lead at $K=1$, it exhibits a rapid performance gain as the support set size increases to $K=3$, establishing a clear dominance thereafter. Crucially, unlike many baselines that trade Recall for Precision, DUPLE achieves the highest performance in both metrics simultaneously. In Figure 5(e), DUPLE’s Precision climbs from $69\%$ at 1-shot to over $84\%$ at 10-shot, significantly outperforming the next best method. This overall advantage is reflected in the F1 score, as shown in Figure 5(f), where DUPLE’s F1 score is more than 10% higher than that of the best baseline model, ProtoNet. Since the F1 Score represents the harmonic mean of Precision and Recall, DUPLE’s top-tier performance on this metric indicates that it provides the most robust solution for practical security monitoring, achieving an optimal balance between false positives and false negatives.
5.2.3 Feature Space Visualization
(a) OSDG1 ($K=5$)
(b) OSDG2 ($K=5$)
Figure 6: t-SNE visualization of the feature embeddings learned by DUPLE (5-shot) on representative held-out deployments. (a) OSDG1(Barbed wire): Smash forms a compact and well-separated cluster, whereas Climb is highly fragmented into multiple scattered sub-clusters and shows noticeable mixing with Background and Rain. This pattern reflects strong deployment-induced variability and helps explain the more challenging performance on OSDG1. (b) OSDG2(Railing):The embeddings exhibit clearer class separation among Background, Climb, and Impact, with substantially reduced inter-class mixing compared with OSDG1, consistent with the stronger quantitative results on OSDG2.
To qualitatively assess the discriminative structure of the learned representations, we employ t-SNE to project the high-dimensional query embeddings into a two-dimensional space for visualization. Under the rigorous Leave-One-Deployment-Out (LODO) protocol, the feature geometry may vary across folds due to deployment-specific signal characteristics. To provide a representative and reproducible view without cherry-picking, we report the embeddings from the most challenging deployment fold for each dataset, identified by the lowest silhouette score in the learned embedding space. Accordingly, the visualizations correspond to the Barbed wire deployment for OSDG1 and the Railing deployment for OSDG2 (Fig. 6).
Table 11: Comparison of computational efficiency.The metrics were measured on a single NVIDIA vGPU-32GB with batch size = 1, using standardized input signals of 150,000 sampling points ( ${\sim}3$ s at 50 kHz). Time (ms) reports the model inference latency excluding STFT preprocessing and data I/O.
Model Params (M) FLOPs (G) Time (ms) FPS
1D-CNN 0.01 0.11 0.13 7690.11
2D-CNN 0.04 0.74 0.15 6840.92
Transformer 1.19 0.31 0.61 1634.76
ProtoNet 2.44 1.65 1.46 685.71
DUPLE (Ours) 2.48 3.30 6.12 163.42
In the more challenging OSDG1 setting, Fig. 6(a) reveals a heterogeneous embedding structure under strong deployment-induced variations. The Smash samples form a compact and well-separated cluster, whereas Climb exhibits pronounced fragmentation into multiple scattered sub-clusters, with noticeable mixing with Background and Rain. This pattern reflects substantial intra-class variability and deployment-specific distribution shifts, which is consistent with the more difficult quantitative results reported in Table 5. Importantly, despite the fragmented structure, the embeddings do not collapse into a single indistinguishable region, indicating that DUPLE still preserves meaningful class-related structures under this challenging deployment.
In contrast, Fig. 6(b) shows that OSDG2 (Railing) yields a more structured organization of the embedding space, with clearer separation among Background, Climb, and Impact and reduced inter-class mixing compared with OSDG1. This observation aligns with the stronger performance of DUPLE on OSDG2, suggesting that the proposed dual-domain and prototype-based metric learning can produce more consistent class-wise clustering under certain deployment conditions.
5.3 Computational Efficiency Analysis
To assess the deployment feasibility of the proposed method in real-time monitoring systems, we benchmarked the model complexity (Parameters), computational cost (FLOPs), and inference latency against representative baselines. All evaluations were conducted on a single NVIDIA VGPU-32GB using a batch size of 1 to simulate the sequential data processing typical of edge computing environments. The standardized input length was set to 150,000 sampling points.
Table 11 summarizes the efficiency metrics. Traditional deep learning models, especially 1D-CNN and 2D-CNN, have the lowest computational overhead, with inference times below 0.2 milliseconds. However, their limited parameter capacities of 0.01 M and 0.04 M, respectively, restrict feature representation and lead to poor generalization performance. While Transformer models maintain a moderate inference speed, their accuracy is lower compared to meta-learning methods, indicating that standard attention mechanisms alone are insufficient for this specific cross-domain few-shot task.
Table 12: Per-Class Accuracy on OSDG1 and OSDG2.
@@ 5.3 Computational Efficiency Analysis
Background Climb Rain Smash Background Climb Impact
XGBoost $24.40\pm 24.40$ $67.04\pm 29.21$ $4.25\pm 3.44$ $38.37\pm 38.37$ $73.59\pm 37.35$ $56.50\pm 41.17$ $45.50\pm 41.32$
SVM $9.60\pm 9.60$ $75.97\pm 14.72$ $43.92\pm 6.08$ $7.86\pm 7.86$ $63.16\pm 38.85$ $64.33\pm 45.58$ $39.17\pm 39.07$
KNN $12.11\pm 4.11$ $3.72\pm 1.28$ $59.24\pm 8.07$ $64.60\pm 34.60$ $68.83\pm 31.03$ $39.06\pm 34.54$ $83.03\pm 17.46$
1DCNN $\pm 0.00 0.00±0.00$$\pm 0.00$ $63.94\pm 36.06$ $76.87\pm 17.41$ $28.67\pm 40.54$ $38.99\pm 37.88$ $87.67\pm 10.00$
2DCNN $43.10\pm 38.30$ $59.46\pm 40.54$ $37.51\pm 30.78$ $41.63\pm 11.23$ $64.32\pm 41.77$ $44.69\pm 39.11$ $66.16\pm 46.78$
LSTM $27.03\pm 27.03$ $32.56\pm 32.56$ $46.15\pm 46.15$ $16.53\pm 14.90$ $54.65\pm 40.19$ $43.92\pm 33.25$ $89.50\pm 13.46$
Transformer $2.00\pm 1.20$ $42.01\pm 19.24$ $25.29\pm 0.29$ $42.43\pm 22.43$ $50.13\pm 8.57$ $70.32\pm 32.78$ $81.78\pm 15.63$
GroupDRO $18.60\pm 18.60$ $58.83\pm 39.92$ $70.13\pm 1.02$ $64.03\pm 31.17$ $98.72\pm 1.81$ $58.96\pm 39.27$ $66.67\pm 47.14$
Coral $11.63\pm 11.63$ $26.20\pm 22.45$ $70.66\pm 28.38$ $90.77\pm 3.63$ $78.52\pm 28.76$ $34.68\pm 45.50$ $100.00\pm 0.00$
ProtoNet $59.25\pm 12.36$ $24.80\pm 24.80$ $37.52\pm 31.43$ $4.62\pm 2.47$ $51.65\pm 33.60$ $73.28\pm 15.32$ $74.56\pm 23.49$
MAML $65.41\pm 29.01$ $61.46\pm 9.45$ $3.29\pm 0.15$ $28.95\pm 14.90$ $56.10\pm 22.56$ $59.41\pm 25.83$ $78.40\pm 23.72$
ANIL $45.13\pm 24.69$ $39.36\pm 4.13$ $15.09\pm 5.73$ $41.77\pm 18.73$ $59.07\pm 38.61$ $61.95\pm 25.35$ $68.91\pm 32.57$
R2D2 $32.14\pm 20.66$ $84.75\pm 14.83$ $31.40\pm 5.66$ $12.68\pm 11.23$ $68.94\pm 29.44$ $58.28\pm 30.93$ $64.79\pm 12.28$
FEAT $23.03\pm 2.10$ $65.06\pm 34.94$ $44.81\pm 26.66$ $66.49\pm 30.57$ $55.59\pm 32.83$ $42.45\pm 42.02$ $75.17\pm 35.11$
DUPLE (Ours) $69.25\pm 29.14$ $64.87\pm 1.56$ $52.13\pm 23.19$ $65.99\pm 17.63$ $71.60\pm 36.82$ $86.41\pm 6.78$ $74.35\pm 19.37$
Compared to the ProtoNet baseline model, our proposed DUPLE framework introduces higher computational load and latency, at 3.30 G FLOPs and 6.12 ms, respectively. This increase is attributed to the parallel processing architecture of the dual-domain branches and the additional computation required for instance-level reweighting by the SGN module. Although DUPLE is slower than lightweight baselines, it still satisfies real-time DFOS monitoring with a large margin (6.12 ms per ${\sim}3$ s window). Nevertheless, DUPLE achieves an inference speed of approximately 163 FPS. Considering that typical distributed fiber optic sensing systems require alarm response frequencies of 20 Hz to 50 Hz, DUPLE’s processing speed is sufficient to support real-time operation. Therefore, the proposed method achieves a practical trade-off, significantly improving accuracy without violating latency constraints in industrial deployments.